The sum of three numbers is 2. If twice the second number is added to the sum of first and third, we get 1. On adding the sum of second and third number to five times the first, we get 6. Find the three number by using matrices.
Answer by stanbon(75887)

The sum of three numbers is 2. If twice the second number is added to the sum of first and third, we get 1. On adding the sum of second and third number to five times the first, we get 6. Find the three number by using matrices.
------
Equation:
x + y + z = 2
2y + x+z = 1
y+z + 5x = 6
----
Rearrange:
x + y + z = 2
x +2y + z = 1
5x+ y + z = 6
----------------
I used the matrix function on my TI-84 to get:
x = 1
y = -1
z = 2
=============
Cheers,
Stan H.
===============